# Effect of Strain on the Band Gap of Monolayer MoS2

### v2

@@ Abstract
Monolayer $\mathrm{MoS_{2}}$ under strain has many interesting properties and possible applications in technology. A recent experimental study examined the effect of strain on the bandgap of monolayer $\mathrm{MoS_{2}}$ on a mildly curved graphite surface, reporting that under biaxial strain with a Poisson’s ratio of 0.44, the bandgap decreases at a rate of 400 meV/strain. In this work, we performed density functional theory (DFT) calculations for a free-standing $\mathrm{MoS_{2}}$ monolayer, using the generalized gradient approximation (GGA) PBE, the hybrid functional HSE06, and many-body perturbation theory with the GW approximation using PBE wavefunctions (G0W0@PBE). We found that under biaxial strain with the experimental Poisson’s ratio, the bandgap decreases at rates of 63 meV/strain (PBE), 73 meV/strain (HSE06), and 43 meV/strain (G0W0@PBE), which are significantly smaller than the experimental rate. We also found that PBE predicts a similarly smaller rate (90 meV/strain) for a different Poisson’s ratio of 0.25. Spin-orbit correction (SOC) has little effect on the gap or its strain dependence. Additionally, we observed a semiconductor-to-metal transition under an equal tensile biaxial strain of 10% and a transition from a direct to an indirect bandgap, consistent with previous theoretical work.
@@ I Introduction
Graphene is a typical two-dimensional (2D) layered material. Monolayer graphene has been found to exhibit several notable features, including high electrical conductivity, high transparency, high thermal conductivity at ambient temperature, a high Young’s modulus, and a high specific surface area Allen et al. (2010); Soldano et al. (2010). Graphene possesses outstanding electrical properties; however, being a zero-gap material Novoselov et al. (2004) limits its application in logical circuits. In an effort to overcome the limitations of graphene and increase its variety of uses, scientists have turned their attention back to alternative 2D materials that resemble graphene Ganatra and Zhang (2014); Chhowalla et al. (2014). Monolayer MoS2 has gained significant interest in recent days.
@@
Monlolayer MoS2 has a 2D nature and a large band gap. Similar to graphene Lee et al. (2008), a single-layer MoS2 has strong inherent tensile strength and flexibility out of plane, allowing the film to withstand strains of up to $11\%$ before rupturing Bertolazzi et al. (2011). According to theoretical predictions, the band gap of semiconducting transition-metal dichalcogenides (TMDCs) would be significantly impacted by tensile strain, and applying approximately 10% biaxial strain might potentially close the gap entirely Johari and Shenoy (2012); Yue et al. (2012); Shi et al. (2013). Strains can be used to control the electronic properties of two-dimensional (2D) materials, which is important for implementation of a 2D material into flexible electronics and next-generation strain engineering devices. Therefore, it becomes important that the material withstands the desired strain, and that we know the effect of the strain on the material. A recent experimental study of an $\mathrm{MoS_{2}}$ monolayer on a mildly curved graphite substrate has measured the relationship between band gap change and strain in monolayer MoS2 Trainer et al. (2019). Trainer et al. Trainer et al. (2019) determined the Poisson’s ratio of monolayer MoS2 to be 0.44. A positive Poisson’s ratio has been previously reported Kang et al. (2015); Woo et al. (2016a, b). Because of a positive Poisson’s ratio, the crystal will respond to a tensile strain in one direction by compressing in the perpendicular direction. Trainer et al. Trainer et al. (2019) found that the quasiparticle band gap decreases at the rate of 400 meV/strain until a nominal strain of 2.5% and at a much slower rate from 2.5% to 4.9%. In this work, we have studied band gap change with strain for a free-standing $\mathrm{MoS_{2}}$ monolayer theoretically, using the HSE06 Krukau et al. (2006) and G0W0 Hedin (1965); Hybertsen and Louie (1985); van Schilfgaarde et al. (2006) methods. We also used the PBE Perdew et al. (1996) method to compare the band gap of an unstrained monolayer MoS2 with the work of Woo et al. Woo et al. (2016a). Our results show that the band gap decreases at a much slower rate than reported in the work by Trainer et al. Trainer et al. (2019).
@@
Figure 1: Rectangular unit cell of monolayer MoS2. Only the S (yellow) atoms in the plane above the plane of the Mo (purple) atoms can be seen. Those in the plane below are obscured by those above. The lattice constant a is the length of the rectangular unit cell in the x-direction and b in the y-direction.
@@ II Computational Methods
We used the rectangular unit cell as shown in Figure 1 for our calculations. The rectangular unit cell helped to strictly maintain a Poisson’s ratio of 0.44. The rectangular unit cell comprises two Mo atoms and four S atoms. The HSE06 and PBE calculations were performed using the Vienna Ab initio Package Simulation (vasp) Kresse and Hafner (1994); Kresse and Joubert (1999). For HSE06 and PBE calculations, we used an 16 x 16 x 2 $\Gamma$-centered k-grid to sample the Brillouin zone. We used a 400 eV cutoff energy for the plane-wave basis for HSE06 and 520 eV for PBE. Initially, we relaxed the lattice parameters a and b and the internal co-ordinates of ions while keeping c fixed at 10 Å to make a three-dimensional super cell. To apply strain, we utilized the relaxed structure and adjusted the lattice parameters a and b to achieve the desired strain, followed by the relaxations of internal co-ordinates of ions. During the strain application, for HSE06 and G0W0@PBE calculations, we maintained the Poisson’s ratio of 0.44 to remain consistent with experiment Trainer et al. (2019). If $\epsilon_{x}$ represents the strain along the x-direction, then the strain along the y-direction ($\epsilon_{y}$) would be $-0.44\epsilon_{x}$. If a and b are the lattice vectors before applying strain, and anew and bnew are new lattice vectors after applying strain, then:
@@
The G0W0 Deslippe et al. (2012); Hybertsen and Louie (1986) calculations were conducted in BerkeleyGW Deslippe et al. (2012) by pairing with Quantum ESPRESSO Giannozzi et al. (2017). The wavefunction energy cutoff is 70 Ry ($\sim$950 eV). The energy cutoff for the epsilon matrix is 18 Ry ($\sim$240 eV). The k-point mesh of 20 × 12 × 1 is used for the rectangular cell. The band number for summation is 220. The correction of the exact static remainder and the 2D slab Coulomb truncation were used.
@@
Table 1: The computed and experimental band gaps (in eV) and changes ($\Delta$) in the band gap relative to the gap at zero strain of monolayer MoS2. The HSE06 and G0W0@PBE results correspond to the rectangular unit cell shown in Fig. 1. Experimental values from Ref. Trainer et al. (2019).
@@
Table 2: band gap vs. strain and lattice parameter (in Å) calculated using PBE for the rectangular cell shown in Fig. 1. Here, we applied uniform biaxial tensile strain.
@@ III Results and Discussion
Figure 2: Change in the band gap $E_{\mathrm{g}}$ as a function of strain $\epsilon_{x}$ along the x-direction using the rectangular unit cell of monolayer MoS2 based on the HSE06 functional. We applied strain such that Poisson’s ratio remains 0.44.
@@
Figure 3: Band structure of (a) unstrained, (b) 1%, (c) 2.5 %, and (d) 4% strained monolayer MoS2 calculated using a rectangular cell with PBE+SOC and G0W0@PBE+SOC methods. We applied strain such that Poisson’s ratio remains 0.44. The plot shows unstrained monolayer MoS2 to be a direct band gap material, and 4% strained monolayer MoS2 to be an indirect band gap material at the PBE level. However, it is still almost direct at the G0W0@PBE level. (By K we indicate the point along the line from $\Gamma$ to X that is closest to the symmetry point K of the hexagonal zone.)
@@
Figure 4: PBE band gap $E_{\mathrm{g}}$ as a function of equal strain $\epsilon$ applied along both the x and y directions using a rectangular unit cell of monolayer MoS2 (equivalent to a Poisson’s ratio of -1).
The band gap of 1H $\mathrm{MoS_{2}}$ monolayer decreases with increasing in-plane tensile strain, a trend opposite to that of some other transition metal dichalcogenides (TMDs), such as the 1T $\mathrm{TiSe_{2}}$ monolayer Pasquier and Yazyev (2018). The decreasing trend in $\mathrm{MoS_{2}}$ can be attributed to the effect on the band centers of the change in ligand field induced by the in-plane strain, the change of the widths of the $d$-orbital bands, and the change of $d$-orbital SOC splitting. A more detailed analysis is presented in the Supplementary Information (SI).
To investigate the effect of spin-orbit correction (SOC), we also calculated the PBE+SOC and G0W0@PBE+SOC band structures of monolayer MoS2 for no strain, 1.0% strain, 2.5% strain, and 4% strain along the x-direction with a strain along the y-direction to maintain Poisson’s ratio of 0.44. The SOC effect is not always significant to band structures of 2D materials. For example, the SOC effect on the bands around the Fermi level of monolayer $\mathrm{Ta_{2}Ni_{3}Te_{5}}$ was shown to be very small Guo et al. (2022). For 1H WSe2 monolayer, the SOC effect induces a strong spin-splitting ($\sim$ 0.5 eV) around the valence band maximum Reyes-Retana and Cervantes-Sodi (2016). Here, for the 1H MoS2 monolayer, the SOC-induced band splitting near the valence band maximum is about 0.15 eV, consistent with a previous study Qiu et al. (2013). As shown in Figure 3a and 3b, both PBE+SOC and G0W0 show the unstrained and 1.0% strained monolayer MoS2 to be a direct band gap semiconductor. At 2.5% strain, PBE+SOC still shows a direct gap, while G0W0 shows that the valence band maximum (VBM) and the conduction band minimum (CBM) located in the $\Gamma$X line and close to X are slightly misaligned, showing a slightly indirect gap. At 4% strain, PBE+SOC shows that the local VBM at Gamma has become the global VBM, yielding an indirect gap. Therefore, with increasing strain this material tends to change from a direct band gap at an indirect band gap material. Such an effect has been previously observed in Refs. Lu et al. (2012); Conley et al. (2013); Johari and Shenoy (2012).
@@ IV Conclusion
In this work, we theoretically studied the band gap and its change in free-standing monolayer MoS2 using the HSE06 and G0W0 methods to compare with the experimental finding for monolayer MoS2 on a mildly curved graphite substrate Trainer et al. (2019). We quantified the rate of decrease of band gap with strain (that maintains Poisson’s ratio of 0.44) for monolayer MoS2. We found the rate of decrease of band gap to be 63 meV/strain for PBE, 73 meV/strain for HSE06 and 45 meV/strain for G0W0@PBE, which are much smaller than the experimental result of 400 meV/strain Trainer et al. (2019). The discrepancy is important to resolve. Furthermore, we demonstrated that PBE and G0W0@PBE predict monolayer MoS2 to be a direct band gap material, with the potential to transition into an indirect band gap material under strain. We also quantified the rate of decrease of band gap with strain (that maintains Poisson’s ratio of 0.25) to be 90 mev/strain at PBE level. Additionally, we observed that 10% equal biaxial tensile strain can transform MoS2 from a semiconductor to a metal at the PBE level, but not at the HSE06 level.
Finally, we briefly comment on the reason why the HSE06 energy gap in the bandstructure of a solid is typically larger and more realistic than the PBE gap. The PBE density functional for the exchange-correlation energy is implemented in the original Kohn-Sham Kohn and Sham (1965) scheme in which the exchange-correlation potential is a function of position that multiplies an orbital, while the HSE06 hybrid of PBE with a fraction of exact exchange depends explicitly on the occupied orbitals and is implemented in a generalized Seidl et al. (1996) Kohn-Sham scheme in which the exchange-correlation potential is a non-local operator. PBE reasonably predicts the bandstructure gap that would be predicted by a position-dependent multiplicative exchange-correlation potential that yields the exact electron density in a solid Perdew and Levy (1983); Sham and Schlüter (1983), while HSE06, to the extent that it correctly predicts ionization energy minus electron affinity for the solid from total energy differences, also predicts the correct fundamental gap from its bandstructure Perdew et al. (2017). The difference between the HSE06 and PBE gaps is an estimate of the discontinuity Perdew and Levy (1983); Sham and Schlüter (1983); Perdew et al. (2017, 1982) of the exact Kohn-Sham potential in an insulating solid as the electron number crosses the number needed to fill a band. Another way to understand this is to note that the electron self-energy that yields the exact quasi-particle spectrum must be a nonlocal operator Sham and Schlüter (1983), like the HSE06 and unlike the PBE and the exact Kohn-Sham exchange-correlation potentials.

### v3

@@ Abstract
Monolayer molybdenum disulfide ($\mathrm{MoS_{2}}$) under strain has many interesting properties and possible applications in technology. A recent experimental study examined the effect of strain on the bandgap of monolayer $\mathrm{MoS_{2}}$ on a mildly curved graphite surface, reporting that under biaxial strain with a Poisson’s ratio of 0.44, the bandgap decreases at a rate of 400 meV/strain. In this work, we performed density functional theory (DFT) calculations for a free-standing $\mathrm{MoS_{2}}$ monolayer, using the generalized gradient approximation (GGA) PBE, the hybrid functional HSE06, and many-body perturbation theory with the GW approximation using PBE wavefunctions (G0W0@PBE). For the unstrained monolayer, we found a standard level of agreement for the bandgap between theory and experiment. For biaxial strain at the experimental Poisson’s ratio, we found that the bandgap decreases at rates of 63 meV/strain (PBE), 73 meV/strain (HSE06), and 43 meV/strain (G0W0@PBE), which are significantly smaller than the experimental rate. We also found that PBE predicts a similarly smaller rate (90 meV/strain) for a different Poisson’s ratio of 0.25. Spin-orbit correction (SOC) has little effect on the gap or its strain dependence. The strong disagreement between theory and experiment may reflect an unexpectedly strong effect of the substrate on the strain dependence of the gap. Additionally, we observed a transition from a direct to an indirect bandgap under strain, and (under an equal biaxial strain of 10%) a semiconductor-to-metal transition, consistent with previous theoretical work.
@@ I Introduction
Graphene is a typical two-dimensional (2D) layered material. Monolayer graphene has been found to exhibit several notable features, including high electrical conductivity, high transparency, high thermal conductivity at ambient temperature, a high Young’s modulus, and a high specific surface area Allen et al. (2010); Soldano et al. (2010). Graphene possesses outstanding electrical properties; however, being a zero-gap material Novoselov et al. (2004) limits its application in logical circuits. In an effort to overcome the limitations of graphene and increase its variety of uses, scientists have turned their attention back to alternative 2D materials that resemble graphene Ganatra and Zhang (2014); Chhowalla et al. (2014). Monolayer MoS2 has gained significant interest over the past decade.
@@
Monlolayer MoS2 has a 2D nature and a large band gap. Similar to graphene Lee et al. (2008), a single-layer MoS2 has strong inherent tensile strength and flexibility out of plane, allowing the film to withstand strains of up to $11\%$ before rupturing Bertolazzi et al. (2011). According to theoretical predictions, the band gap of semiconducting transition-metal dichalcogenides (TMDCs) would be significantly impacted by tensile strain, and applying approximately 10% biaxial strain might potentially close the gap entirely Johari and Shenoy (2012); Yue et al. (2012); Shi et al. (2013). Strains can be used to control the electronic properties of two-dimensional (2D) materials, which is important for implementation of a 2D material into flexible electronics and next-generation strain engineering devices. Therefore, it becomes important that the material withstands the desired strain, and that we know the effect of the strain on the material. A recent experimental study of an $\mathrm{MoS_{2}}$ monolayer on a macroscopically curved surface (with a radius of curvature greater than 107 Å) has measured a relationship between band gap change and strain in monolayer MoS2 Trainer et al. (2019). Trainer et al. Trainer et al. (2019) determined the Poisson’s ratio of monolayer MoS2 to be 0.44. A positive Poisson’s ratio has been previously reported Kang et al. (2015); Woo et al. (2016a). Because of a positive Poisson’s ratio, the crystal will respond to a tensile strain in one direction by compressing in the perpendicular direction. Trainer et al. Trainer et al. (2019) applied a tensile strain along an in-plane direction and found that the quasiparticle band gap decreases at the rate of 400 meV/strain until a nominal strain of 2.5% and at a much slower rate from 2.5% to 4.9%. In this work, we have studied band gap change with strain for a free-standing $\mathrm{MoS_{2}}$ monolayer theoretically, using the Heyd-Scuseria-Ernzerhof screened hybrid functional HSE06 Krukau et al. (2006) and G0W0 Hedin (1965); Hybertsen and Louie (1985); van Schilfgaarde et al. (2006) methods. We also used the Perdew-Burke-Ernzerhof (PBE) Perdew et al. (1996) method to compare the band gap of an unstrained monolayer MoS2 with the work of Woo et al. Woo et al. (2016a). Our results show that the band gap decreases at a much slower rate than reported in the work by Trainer et al. Trainer et al. (2019).
@@
Figure 1: Orthorhombic unit cell of monolayer MoS2. Only the S (yellow) atoms in the plane above the plane of the Mo (purple) atoms can be seen. Those in the plane below are obscured by those above. The lattice constant a is the length of the orthorhombic unit cell in the x-direction and b in the y-direction.
@@ II Computational Methods
We used the orthorhombic unit cell as shown in Figure 1 for our calculations. The orthorhombic unit cell helped to strictly maintain a Poisson’s ratio of 0.44. The orthorhombic unit cell comprises two Mo atoms and four S atoms. The HSE06 and PBE calculations were performed using the Vienna Ab initio Simulation Package (vasp) Kresse and Hafner (1994); Kresse and Joubert (1999). For HSE06 and PBE calculations, we used a 16 x 16 x 2 $\Gamma$-centered k-grid to sample the Brillouin zone. We used a 400 eV cutoff energy for the plane-wave basis for HSE06 and 520 eV for PBE. Initially, we relaxed the lattice parameters a and b and the internal co-ordinates of ions while keeping c fixed at 10 Å to make a three-dimensional super cell. To apply strain, we utilized the relaxed structure and adjusted the lattice parameters a and b to achieve the desired strain, followed by the relaxations of internal co-ordinates of ions. During the strain application, for HSE06 and G0W0@PBE calculations, we maintained the Poisson’s ratio of 0.44 to remain consistent with experiment Trainer et al. (2019). If $\epsilon_{x}$ represents the strain along the x-direction, then the strain along the y-direction ($\epsilon_{y}$) would be $-0.44\epsilon_{x}$. If a and b are the lattice vectors before applying strain, and anew and bnew are new lattice vectors after applying strain, then:
@@
The G0W0 Deslippe et al. (2012); Hybertsen and Louie (1986) calculations were conducted in BERKELEYGW Deslippe et al. (2012) by pairing with QUANTUM ESPRESSO (QE) Giannozzi et al. (2017). The wavefunction energy cutoff is 70 Ry ($\sim$950 eV). The energy cutoff for the dielectric matrix ($\epsilon$) is 18 Ry ($\sim$240 eV). A k-point mesh of 20 × 12 × 1 is used for the orthorhombic cell. The band number for summation is 220. The correction of the exact static remainder and the 2D slab Coulomb truncation were used. The pseudopotentials used in QE are ONCVPSP (Optimized Norm-Conserving Vanderbilt Pseudopotential) fully-relativistic version. For atom Mo, the valence electrons are 4s2/4p6/4d5/5s1. For atom S, the valence electrons are 3s2/3p4.
The different k-meshes used in the different codes are both essentially converged for the band gap. There would be no difference between one and two k-points in the direction out of the plane if the out-of-plane cell dimension were large enough. Increasing this cell dimension from 10 Å to 15 Å in VASP increased the PBE gap by only 0.05 eV.
@@
Table 1: The computed and experimental band gaps (in eV) and changes ($\Delta$) in the band gap relative to the gap at zero strain of monolayer MoS2. The HSE06 and G0W0@PBE results correspond to the orthorhombic unit cell shown in Fig. 1. Experimental values from Ref. Trainer et al. (2019).
@@
Table 2: band gap vs. strain and lattice parameter (in Å) calculated using PBE for the orthorhombic cell shown in Fig. 1. Here, we applied uniform biaxial tensile strain.
@@ III Results and Discussion
Figure 2: Change in the band gap $E_{\mathrm{g}}$ as a function of strain $\epsilon_{x}$ along the x-direction using the orthorhombic unit cell of monolayer MoS2 based on the HSE06 functional. We applied strain such that Poisson’s ratio remains 0.44.
@@
Figure 3: Band structure of (a) unstrained, (b) 1%, (c) 2.5 %, and (d) 4% strained monolayer MoS2 calculated using a orthorhombic cell with PBE+SOC and G0W0@PBE+SOC methods. We applied strain such that Poisson’s ratio remains 0.44. The plot shows unstrained monolayer MoS2 to be a direct band gap material, and 4% strained monolayer MoS2 to be an indirect band gap material at the PBE level. However, it is still almost direct at the G0W0@PBE level. (By K we indicate the point along the line from $\Gamma$ to X that is closest to the symmetry point K of the hexagonal zone.)
@@
Figure 4: PBE band gap $E_{\mathrm{g}}$ as a function of equal strain $\epsilon$ applied along both the x and y directions using a orthorhombic unit cell of monolayer MoS2 (equivalent to a Poisson’s ratio of -1).
The band gap of 1H $\mathrm{MoS_{2}}$ monolayer decreases with increasing in-plane tensile strain, a trend opposite to that of some other transition metal dichalcogenides (TMDs), such as the 1T $\mathrm{TiSe_{2}}$ monolayer Pasquier and Yazyev (2018). The decreasing trend in $\mathrm{MoS_{2}}$ can be attributed to the effect on the band centers of the change in ligand field induced by the in-plane strain, the change of the widths of the $d$-orbital bands, and the change of $d$-orbital SOC splitting. A more detailed analysis is presented in the Supplementary Information (SI) SM2024 , which includes Refs. Chhowalla et al. (2014); Tang et al. (2023); Wilson and Yoffe (1969).
To investigate the effect of spin-orbit correction (SOC), we also calculated the PBE+SOC and G0W0@PBE+SOC band structures of monolayer MoS2 for no strain, 1.0% strain, 2.5% strain, and 4% strain along the x-direction with a strain along the y-direction to maintain Poisson’s ratio of 0.44. The SOC effect is not always significant for band structures of 2D materials. For example, the SOC effect on the bands around the Fermi level of monolayer $\mathrm{Ta_{2}Ni_{3}Te_{5}}$ was shown to be very small Guo et al. (2022). For 1H WSe2 monolayer, the SOC effect induces a strong spin-splitting ($\sim$ 0.5 eV) around the valence band maximum Reyes-Retana and Cervantes-Sodi (2016). Here, for the 1H MoS2 monolayer, the SOC-induced band splitting near the valence band maximum is about 0.15 eV, consistent with a previous study Qiu et al. (2013). As shown in Figure 3a and 3b, both PBE+SOC and G0W0 show the unstrained and 1.0% strained monolayer MoS2 to be a direct band gap semiconductor. At 2.5% strain, PBE+SOC still shows a direct gap, while G0W0 shows that the valence band maximum (VBM) and the conduction band minimum (CBM) located in the $\Gamma$X line and close to X are slightly misaligned, showing a slightly indirect gap. At 4% strain, PBE+SOC shows that the local VBM at $\Gamma$ has become the global VBM, yielding an indirect gap. Therefore, with increasing strain this material tends to change from a direct band gap at an indirect band gap material. Such an effect has been previously observed in Refs. Lu et al. (2012); Conley et al. (2013); Johari and Shenoy (2012).
@@ IV Conclusion
In this work, we theoretically studied the band gap and its change in free-standing monolayer MoS2 using the HSE06 and G0W0 methods to compare with the experimental finding for monolayer MoS2 on a mildly curved graphite substrate Trainer et al. (2019). We quantified the rate of decrease of band gap with strain (that maintains Poisson’s ratio of 0.44) for monolayer MoS2. We found the rate of decrease of band gap to be 63 meV/strain for PBE, 73 meV/strain for HSE06 and 45 meV/strain for G0W0@PBE, which are much smaller than the experimental result of 400 meV/strain Trainer et al. (2019). The discrepancy is important to resolve. It suggests the possibility of an unexpectedly strong effect of the substrate on the strain dependence of the gap. Furthermore, we demonstrated that PBE and G0W0@PBE predict monolayer MoS2 to be a direct band gap material, with the potential to transition into an indirect band gap material under strain. We also quantified the rate of decrease of band gap with strain (that maintains Poisson’s ratio of 0.25) to be 90 mev/strain at PBE level. Additionally, we observed that 10% equal biaxial tensile strain can transform MoS2 from a semiconductor to a metal at the PBE level, but not at the HSE06 level.
Finally, we briefly comment on the reason why the HSE06 energy gap in the bandstructure of a solid is typically larger and more realistic than the PBE gap. The PBE density functional for the exchange-correlation energy is implemented in the original Kohn-Sham Kohn and Sham (1965) scheme in which the exchange-correlation potential is a function of position that multiplies an orbital, while the HSE06 hybrid of PBE with a fraction of exact exchange depends explicitly on the occupied orbitals and is implemented in a generalized Seidl et al. (1996) Kohn-Sham scheme in which the exchange-correlation potential is a non-local operator. PBE reasonably predicts the bandstructure gap that would be predicted by a position-dependent multiplicative exchange-correlation potential that yields the exact electron density in a solid Perdew and Levy (1983); Sham and Schlüter (1983), while HSE06, to the extent that it correctly predicts ionization energy minus electron affinity for the solid from total energy differences, also predicts the correct fundamental gap from its bandstructure Perdew et al. (2017). The difference between the HSE06 and PBE gaps is an estimate of the discontinuity Perdew and Levy (1983); Sham and Schlüter (1983); Perdew et al. (2017, 1982) of the exact Kohn-Sham potential in an insulating solid as the electron number crosses the number needed to fill a band. Another way to understand this is to note that the electron self-energy that yields the exact quasi-particle spectrum must be a nonlocal operator Sham and Schlüter (1983), like HSE06 and unlike the PBE and the exact Kohn-Sham exchange-correlation potentials.